# Evidence of equilibrium dynamics in human social networks evolving in time

### v2

@@ Discussion
In this paper, we have shown that the dynamical process driving the evolution of the personal relationships network of students belonging to a high school in Madrid, from 2020 to 2024, is stationary, and that the network itself is in dynamical equilibrium. We find that the abundances of edges, triangles and degrees are statistically equivalent in all snapshots and can be accurately predicted by the average transition matrix. Furthermore, a key feature of equilibrium in the statistical mechanics sense, Detailed Balance, is also verified with very small, point deviations.
To the best of our knowledge, this result goes much beyond what has been studied so far. Some previous research reported a certain stability in the data analysed [11, 14, 8, 16, 20, 19, 43, 41], but they neither explored this stability further nor showed rigorously that an empirical network is in equilibrium. In addition, as we mentioned in the introduction, other studies have found stability and robustness in social behaviour in broader populations. For instance, in [34], they show that social networks exhibit a resilience mechanism after the death of a friend, recovering the same number of active connections over time through increased interactions between friends of the decedent. This would be an example of how the network evolves to maintain its structural integrity. They also show that interactions between friends stabilize after a year following the death of a friend, similar to the stationarity in the transition probabilities in our findings. Moreover, they mention the possibility of the existence of a lower bound on individuals’ level of social connection, that would force them to replace lost friendships more quickly than they are driven to establish friendships in general. Within the context of mobility patterns, in [35], the authors found that the number of locations an individual visits regularly is conserved over time, even while individual routines are unstable in the long term because of the continuous exploration of new locations. Furthermore, they find a connection between this number and the the Dunbar’s number, establishing a relation between these mobility patterns and the maintenance of their social relations. In [36, 37, 38] the authors explore how online social behaviour follows consistent patterns over time, with [37] focusing on the role played by ties of different strengths, and proving how different dynamical mechanisms compete for this stability to arise, an idea we will develop on the following discussion. Finally, in [44] the authors show how different group social behaviours are consistent within both students and adult groups, pointing towards the generality in the dynamical behaviour of social bonds.
@@
We now argue that, within this theoretical framework, the concept of equilibrium arises naturally. Basically, in the evolution of personal relationships, the context, the institutional setting, all influence with whom we interact (and, hopefully, with whom we establish relationships). Once this context is stable, people establish relationships allocating their cognitive and material resources among the people with whom they share this social environment. In the broader picture of the network, the other dynamical mechanisms shape how these relationships are structured (in simple terms, a person allocates some resources to establish their relationships, and dynamical mechanisms determine who these people are within the broader network). Thus, every time a relationship disappears, there is a liberation of resources that can be reinvested in new relationships. This argument is supported by the findings we mentioned from [34]. At the level of dynamical mechanisms, there is a balancing process as well. For instance, when this new relationship is established, the reciprocation mechanisms will force it to be reciprocal. Nonetheless, there are other mechanisms destroying this reciprocity, like the formation of hierarchical structures. Similarly, there is a tendency towards the creation of balanced triangles, but also a tendency to avoid conflict and a high random component in conflict, that promotes the destruction of these triangles and the creation of unbalanced ones. Therefore, there is a competition between mechanisms having opposing effects on the network structures. The equilibrium is the expected outcome of the network dynamics.
@@
If indeed social networks are generally in equilibrium, the findings of studies using cross-sectional data could potentially be generalized to the entire unobserved evolution of the network. When studied at the proper temporal and spatial scales after the transient, the observation of single snapshots can provide robust information about the dynamical mechanisms driving the network evolution and the structural features observed. This provides solid grounds for using methods like Exponential Random Graph Models and other methods that compare the network observation to null-models to understand the structural trends observed. From the social sciences perspective, this finding can avoid the difficulties in collecting longitudinal data on social networks, reducing the burden on respondents. Also, it opens the door for the design of intervention studies. Since the network properties are stable over time, the researcher can introduce interventions and associate the observed changes to the intervention isolated from other potential drivers of the network evolution.
@@ 3.1 Materials
The collection of our data was performed through surveys administered in the school via a computer interface. To elicit relationships, students were presented with a list of all other students in the high school. They were then asked to select individuals with whom they had a relationship. Specifically, the questionnaire included the following question: ‘You can now see the list of all the students in the school. Please mark those you have any relationship with by clicking ‘very good relationship’, ‘good relationship’, ‘bad relationship’ or ‘very bad relationship’. Only one choice is possible. If you do not mark any option, it will be understood to mean that you do not have a relationship with the person’. Typically, it took students about 15 minutes to complete the survey, and they were supervised by a school teacher throughout the process. Our Institutional Review Boards (IRB) stipulated an opt-out procedure. We should note that there were no opt-outs, effectively eliminating any potential selection bias. The only students who did not participate were those who were absent on the day of the experiment.
@@
All the aggregated data necessary to replicate our results can be found in this repository.
@@ 3.2 Methods
Let us introduce the concept of a transition matrix using the edges as our macroscopic property. In the case of edges, there are 25 possible edge states. If we take two consecutive snapshots of the network, for each individual edge, we can record what kind of transition it went through. A +2+2 edge can stay a +2+2, or become a +2+1, or even a +2-1, etc. If we repeat this process for all the links, we can construct a 25x25 matrix in which each row is the edge state before the transition, in the first snapshot, and each column is the edge state after the transition, in the second snapshot. The $ij$ element of this matrix would be the number of edges starting in state $i$ and ending in state $j$. In this matrix, we can divide each element by the total count in each row. By doing this, the $ij$ element in this final matrix represents $P(j|i)$, i. e., the conditional probability of an edge of going to the $j$ state provided it started in the $i$ state. We call this matrix the transition matrix $P_{\beta\alpha}$ (also known as the Markov Matrix), from snapshot $\alpha$ to snapshot $\beta$.
@@
We want to analyse if the nine transition matrices are statistically equivalent between them. To simplify the process, we check whether the nine transition matrices are statistically equivalent to the average transition matrix computed by averaging the nine individual matrices. To compare each transition matrix with the average, we resampled the individual matrix randomly from the data 1000 times. We calculated the distance between each element of the randomly sampled matrix and the corresponding element in the average transition matrix. To do such resampling, we created a pool with all the observed transitions for each individual edge in each pair of waves, removing the information about the actual waves in which that individual transition occurred. We then created each randomly resampled matrix by drawing samples at random from this pool with no information about the specific waves in which transitions occurred. We counted how often each element of the resampled matrix was farther from the corresponding element in the average matrix than the actual value for the matrix we wanted to compare. In summary, we calculated, element by element, the probability that a randomly sampled matrix element would be farther from the corresponding element of the average matrix than the same element in the actual transition matrix we were analysing.
It is worth noting that we perform an element-wise comparison. Although we could compare the matrices themselves using metrics like the Frobenius distance, we chose a granular approach to be able to identify individual differences. However, it is important to take into account that this method involves multiple comparisons of elements in a sampled matrix; specifically, we are comparing 625 elements against another 625 elements in the case of the edges transition matrix. Consequently, we expect some elements to deviate more than expected by chance, even if the matrices are equivalent. For such cases, we compute a z-score to determine how significant these deviations are from randomness.
@@ 4 Acknowledgements
M.A.G.-C. acknowledges support from the Comunidad de Madrid through the grants for the hiring of pre-doctoral research personnel in training (reference PIPF-2023/COM-29487). M.A.G.-C. and A.S. acknowledge support from grant PID2022-141802NB-I00 (BASIC) funded by MCIN/AEI/10.13039/501100011033 and by ‘ERDFA way of making Europe’. A. S. T acknowledges support by FCT – Fundação para a Ciência e Tecnologia – through the LASIGE Research Unit, ref. UIDB/00408/2020 (https://doi.org/10.54499/UIDB/00408/2020) and ref. UIDP/00408/2020 (10.54499/UIDP/00408/2020)

### v4

@@ Discussion
In this paper, we have shown that the dynamical process driving the evolution of the personal relationships network of students belonging to a high school in Madrid, from 2020 to 2024, is stationary, and that the network itself is in dynamical equilibrium. We find that the abundances of edges, triangles, degrees and even a more complex metric like the Betweenness Centrality, are statistically equivalent in all snapshots and can be accurately predicted by the average transition matrix. Furthermore, a key feature of equilibrium in the statistical mechanics sense, Detailed Balance, is also verified with very small, point deviations.
To the best of our knowledge, this result goes much beyond what has been studied so far. Some previous research reported a certain stability in the data analysed [11, 14, 8, 16, 20, 19, 47, 45], but they neither explored this stability further nor showed rigorously that an empirical network is in equilibrium. In addition, as we mentioned in the introduction, other studies have found stability and robustness in social behaviour in broader populations. For instance, in [34], they show that social networks exhibit a resilience mechanism after the death of a friend, recovering the same number of active connections over time through increased interactions between friends of the decedent. This would be an example of how the network evolves to maintain its structural integrity. They also show that interactions between friends stabilize after a year following the death of a friend, similar to the stationarity in the transition probabilities in our findings. Moreover, they mention the possibility of the existence of a lower bound on individuals’ level of social connection, that would force them to replace lost friendships more quickly than they are driven to establish friendships in general. Within the context of mobility patterns, in [35], the authors found that the number of locations an individual visits regularly is conserved over time, even while individual routines are unstable in the long term because of the continuous exploration of new locations. Furthermore, they find a connection between this number and the Dunbar’s number, establishing a relation between these mobility patterns and the maintenance of their social relations. In [36, 37, 38] the authors explore how online social behaviour follows consistent patterns over time, with [37] focusing on the role played by ties of different strengths, and proving how different dynamical mechanisms compete for this stability to arise, an idea we will develop on the following discussion. Finally, in [48] the authors show how different group social behaviours are consistent within both students and adult groups, pointing towards the generality in the dynamical behaviour of social bonds.
@@
We now argue that, within this theoretical framework, the concept of equilibrium arises naturally. Basically, in the evolution of personal relationships, the context, the institutional setting, all influence with whom we interact (and, hopefully, with whom we establish relationships). Once this context is stable, people establish relationships allocating their cognitive and material resources among the people with whom they share this social environment. In the broader picture of the network, the other dynamical mechanisms shape how these relationships are structured (in simple terms, a person allocates some resources to establish their relationships, and dynamical mechanisms determine who these people are within the broader network). Thus, every time a relationship disappears, there is a liberation of resources that can be reinvested in new relationships. This argument is supported by the findings we mentioned from [34]. At the level of dynamical mechanisms, there is a balancing process as well. For instance, when this new relationship is established, the reciprocation mechanisms will force it to be reciprocal. Nonetheless, there are other mechanisms destroying this reciprocity, like the formation of hierarchical structures. Similarly, there is a tendency towards the creation of balanced triangles, but also a tendency to avoid conflict and a high random component in conflict, that promotes the destruction of these triangles and the creation of unbalanced ones. Therefore, there is a competition between mechanisms having opposing effects on the network structures. The equilibrium is the expected outcome of the network dynamics. It is worth mentioning that modelling the evolution of the network by integrating these spectrum of mechanisms is a task we are addressing in a future study, trying to explicitly tests all hypotheses presented here.
@@
Although we propose that our findings may generalize to broader populations, we do not claim that every specific result presented here will replicate exactly in other contexts. We acknowledge that our study is based on a specific sample—a high school network representing a particular demographic group within a distinct cultural and institutional setting. Our main argument is that equilibrium behaviour emerges from a balance between competing mechanisms and between necessities and resources allocation. However, the precise operation of these mechanisms may vary across different sample groups. In practice, this implies that one might observe different equilibrium distributions, with variations in the abundances of degrees, edges, and triangles, and differences in the transition matrices. For example, as adolescents mature, they may develop improved emotional regulation and conflict management skills, resulting in fewer conflicts and reduced volatility. Nonetheless, we believe that even in such settings, the system would still reach an equilibrium state driven by the interplay of these mechanisms, albeit in a context-dependent manner.
@@
It is important to recognize that in the work presented here we assume that the system is Markovian, meaning that each subsequent network snapshot depends solely on the current snapshot, with no influence from past states. In other words, we assume that the system has no memory. Nonetheless, we acknowledge that the network snapshots are likely not entirely independent, and that previous positive and negative links may influence future connections. For example, a new positive relationship might emerge with a higher probability between two individuals who previously shared a positive connection, or with a lower probability between those who had a conflict. While our results demonstrate remarkable accuracy in predicting the stationary state under the Markovian assumption—and we do not expect memory effects to significantly alter the overall distributions and aggregated transition matrices—we believe that such effects may still impact the network’s evolution. Specifically, memory effects could modify the probability of certain transitions based on past interactions, so that even if the aggregated probabilities remain unchanged, individual edges might exhibit more variability in their activity over time. This would lead to a heterogeneity in transition probabilities that is not captured by the Markovian approximation. We recognize that this topic deserves further investigation, and we plan to develop a more comprehensive analysis of memory effects and non-Markovian dynamics in future work.
In any case, if indeed social networks are generally in equilibrium, the findings of studies using cross-sectional data could potentially be generalized to the entire unobserved evolution of the network. When studied at the proper temporal and spatial scales after the transient, the observation of single snapshots can provide robust information about the dynamical mechanisms driving the network evolution and the structural features observed. This provides solid grounds for using methods like Exponential Random Graph Models and other methods that compare the network observation to null-models to understand the structural trends observed. From the social sciences perspective, this finding can avoid the difficulties in collecting longitudinal data on social networks, reducing the burden on respondents. Also, it opens the door for the design of intervention studies. Since the network properties are stable over time, the researcher can introduce interventions and associate the observed changes to the intervention isolated from other potential drivers of the network evolution. For example, consider an integration intervention aimed at promoting the inclusion of individuals at the lower end of the positive in-degree distribution, enhancing their visibility and likeability within the network. The impact of this intervention could be evaluated by comparing the resulting degree distributions with the previously observed equilibrium distributions. Similarly, an intervention focused on improving emotional regulation and relationship management might promote more balanced interactions within the network. By analyzing the edge-type distributions before and after such an intervention, one could determine whether the frequency of imbalanced relationships (e.g., +2+1, +1+0, and +2+0) decreases while balanced positive interactions (e.g., +1+1 and +2+2) become more prevalent.
@@ 3.1 Materials
The collection of our data was performed through surveys administered in the school via a computer interface. To elicit relationships, students were presented with a list of all other students in the high school. They were then asked to select individuals with whom they had a relationship. Specifically, the questionnaire included the following question: ‘You can now see the list of all the students in the school. Please mark those you have any relationship with by clicking ‘very good relationship’, ‘good relationship’, ‘bad relationship’ or ‘very bad relationship’. Only one choice is possible. If you do not mark any option, it will be understood to mean that you do not have a relationship with the person’. Typically, it took students about 15 minutes to complete the survey, and they were supervised by a school teacher throughout the process. Our study was approved by the Institutional Review Board of the UC3M, which stipulated an opt-out procedure. We should note that there were no opt-outs, effectively eliminating any potential selection bias. The only students who did not participate were those who were absent on the day of the experiment.
@@
All the aggregated data necessary to replicate our results can be found in this repository. Besides, in the same repository one can find all the codes necessary to reproduce our computations.
@@ 3.2 Methods
Missing Data Treatment
In reconstructing the network snapshots from the survey data, a link from node i to node j can be absent for two distinct reasons. First, it may indicate that person i deliberately chose not to declare any relationship—positive or negative—with person j. Alternatively, the absence of a link may result from person i being absent on the day of the survey, so that the link is missing simply because no response was recorded. In our analysis, we have carefully distinguished between these two cases. Throughout the paper, we use ‘+0’ to denote an absent link that results from a deliberate decision by the respondent not to assign any weight. This is distinct from ‘missing links’, which refer to potential links that are absent due to non-response. Missing links are excluded from all counts to avoid introducing bias. For example, a notation such as ‘+0+1’ indicates that the first person deliberately chose not to reciprocate the link, whereas if the data were missing, that particular link would not be counted in the frequency statistics. All missing data are systematically removed from our computations (including those for stationarity matrices, detailed balance, abundance measures, etc.). For instance, if there is a missing link from i to j and a +1 link from j to i—which would form an edge noted as (miss)+1—this case is not counted as a +0+1 edge. Similarly, if in the next wave both individuals are present and we observe a +1+1 link, this is not regarded as a transition from a +0+1 to a +1+1 state.
@@
Let us introduce the concept of a transition matrix using the edges as our macroscopic property. In the case of edges, there are 25 possible edge states. If we take two consecutive snapshots of the network, for each individual edge, we can record what kind of transition it went through. A +2+2 edge can stay a +2+2, or become a +2+1, or even a +2-1, etc. If we repeat this process for all the links, we can construct a 25x25 matrix in which each row is the edge state before the transition, in the first snapshot, and each column is the edge state after the transition, in the second snapshot. The $ij$ element of this matrix would be the number of edges starting in state $i$ and ending in state $j$. In this matrix, we can divide each element by the total count in each row. By doing this, the $ij$ element in this final matrix represents $P(j|i)$, i. e., the conditional probability of an edge of going to the $j$ state provided it started in the $i$ state. We call this matrix the transition matrix (also known as the Markov Matrix).
@@
Figure 5: Illustration of the method used to construct and compare the transition matrices. To simplify the explanation, an example with three snapshots of a directed network of fixed composition is shown, where links take a value of either 0 or 1.
The purpose is to analyse whether the nine transition matrices are statistically equivalent to one another. However, to simplify the process, we instead check whether each of the nine transition matrices is statistically equivalent to the average transition matrix computed from all the transitions observed between consecutive snapshots. Let us explain the method to perform this statistical comparison. Besides, in Figure 5 we have included a diagram with the main points of the procedure.
Let us introduce some notation using the dynamics of edges as an example. To construct each transition matrix, we record all transitions between edge states that occur from one snapshot to the next. Denote a transition by the letter $K=A,B,C,\dots$, and define $N^{K}$ as the number of transitions observed between one snapshot and the next, for that particular case. For instance, $N^{A}$ is the number of transitions recorded from snapshot 0 to snapshot 1, which we use to construct the transition matrix $M^{A}$ (containing the transition counts as defined in the previous section) and $m^{A}$ (containing the corresponding conditional probabilities, i.e., $P(j|i)$, as defined previously). In summary, $M_{ij}^{A}$ represents the number of transitions observed from edge type $i$ to edge type $j$ between snapshot 0 and snapshot 1, and $m_{ij}^{A}$ represents the conditional probability $P(j|i)$ for that interval. We construct one such matrix for each pair of consecutive snapshots.
By aggregating the observed transitions from all pairs of consecutive snapshots, we form a pool of transitions. This pool contains $N^{A}$ transitions from the first pair, $N^{B}$ transitions from the second pair, and so on, so that the total number of transitions is $N=N^{A}+N^{B}+N^{C}+\cdots$ By aggregating all these transitions, we can construct the Average Transition Matrix $\langle m\rangle$ for the entire observed process.
With this setup in mind, our objective is to determine whether $m^{K}$ (for $K=A,B,C,\dots$) is equivalent to $\langle m\rangle$; that is, whether each individual transition matrix is statistically equivalent to the average matrix. The key idea is that if the process is stationary, each transition matrix between consecutive snapshots is simply a resampling of the average matrix because the underlying process remains constant over time. In other words, we ask whether $m^{K}$ can be interpreted as a resampling of $\langle m\rangle$.
The procedure is as follows. The matrix $m^{K}$, which we wish to compare with $\langle m\rangle$, is constructed from the $N^{K}$ observed transitions between consecutive snapshots. To assess its equivalence to $\langle m\rangle$, we compare it with resampled versions of the average matrix. Specifically, by randomly selecting $N^{K}$ transitions from the pool of all observed transitions, we effectively construct a transition matrix that is, by definition, statistically equivalent to $\langle m\rangle$, but computed with the same number of transitions as $m^{K}$. We generate 1000 such matrices, denoted $\langle m\rangle^{R1}$, $\langle m\rangle^{R2}$, $\dots$, $\langle m\rangle^{R1000}$, to represent random samples statistically equivalent to the average matrix. Our goal is to determine whether these resampled matrices behave similarly to $m^{K}$.
For each element $ij$ of the matrices, we compute the distance between the resampled matrices and the average matrix:
$D_{ij}^{Rx}=\langle m_{ij}\rangle^{Rx}-\langle m_{ij}\rangle\quad(x=1,2,\dots,1000).$
These distances form what we call the ‘distribution of null distances’, which represents the expected differences between $\langle m_{ij}\rangle$ and the corresponding element of a statistically equivalent matrix resampled with $N^{K}$ transitions.
Next, we compare the observed distance
$D_{ij}^{K}=m_{ij}^{K}-\langle m_{ij}\rangle,$
with this null distribution. If $D_{ij}^{K}$ is compatible with the distribution of null distances, then for element $ij$ the real transition matrix $m^{K}$ behaves like a resample of $\langle m\rangle$, and they can be considered statistically equivalent. Note that for each element $ij$ we obtain a distribution of null distances against which the observed distance is compared.
From the distribution of null distances, we compute the following quantities for each element $ij$ of each transition matrix:
  • p-value: Computed as the proportion of times the absolute observed distance is smaller than the absolute null distance, i.e.,
    $\Pr\Bigl{(}|D_{ij}^{K}|<|D_{ij}^{Rx}|\Bigr{)}.$
  • z-score: Computed as
    $z=\frac{D_{ij}^{K}-\mu(D_{ij}^{Rx})}{\sigma(D_{ij}^{Rx})},$
    where $\mu(D_{ij}^{Rx})$ and $\sigma(D_{ij}^{Rx})$ are the mean and standard deviation of the null distance distribution, respectively. This z-score indicates how many standard deviations the observed distance deviates from the mean of the null distribution.
In principle, choosing a significance level of 0.05, we would consider $m_{ij}^{K}$ statistically equivalent to $\langle m_{ij}\rangle$ if the p-value exceeds 0.05. However, because this method involves multiple comparisons across the elements of the matrix, we apply a Bonferroni correction and consider the two elements statistically equivalent only if the p-value exceeds (0.05/number of comparisons). For elements with p-values below this threshold, we analyze the computed z-scores to assess how far the observed element $ij$ deviates from the null distribution.
Estimation of confidence intervals
In our analysis, whenever we need to estimate the confidence interval associated with a probability or a density, computed as $p=x/M$, where $x$ is the number of cases that belong to the category of interest and $M$ is the total number of cases, we use the approximation to the variance of the Binomial distribution such that a 0.95 confidence interval is given by:
$CI(0.95)\approx p\pm 1.96\sqrt{\frac{p(1-p)}{M}}.$
All confidence intervals presented in the paper correspond to 95% confidence intervals.
@@ 4 Acknowledgements
M.A.G.-C. acknowledges support from the Comunidad de Madrid through the grants for the hiring of pre-doctoral research personnel in training (reference PIPF-2023/COM-29487). M.A.G.-C. and A.S. acknowledge support from grant PID2022-141802NB-I00 (BASIC) funded by MCIN/AEI/10.13039/501100011033 and by ‘ERDFA way of making Europe’, and from grant MapCDPerNets—Programa Fundamentos de la Fundación BBVA 2022. A. S. T acknowledges support by FCT – Fundação para a Ciência e Tecnologia – through the LASIGE Research Unit, ref. UID/000408/2025.